6. For a certain function f(x) we have: f(x) = (x - 3)²(2x - 3) and • f(x) = 6(x - 3)(x - 2) (a) Use f to find the interva
Answer #1

Quey 4/2) = (2+ 37² (2x-3) {() = 662-3)(2-2) Anicos Analysic f(x) = (x-33262x<3) 42-33320 I, (2x-3) 20 p=2 x 3 & x= 3 are c(قل) لم مملحه Internal of Concave up = (-09,2) U13,0) Interval of Concave Down = (2,3) (4) kausing pan (4.4) - 143) f (x) = (

Please appreciate my work.
If you have any query regarding this question. You are free to ask.